nick is twice as old as Noah. Nine years ago, nick was five times as old as Noah. How old is each now?
Answer by mananth(16949)

Noah's age =x
Nick's age =2x
9 years ago
Noah =x-9
Nick=2x-9
2x-9=5(x-9)
2x-9 =5x-45
2x-5x=-45+9
-3x=-36
/-3
x=12 noah's age
Nick's age = 12*12 = 24 years